OpenStudy (anonymous):

Aldonza needs to fence in a corral with area 600 square feet (see illustration below). One side (one of the x sides) needs a stronger fence to keep the cows out--materials will cost $8 per foot for this side. The other three sides can be fenced with less expensive materials at $4 per foot. Write an equation using x and y to express Aldonza's cost to build this rectangular fence: Cost =

OpenStudy (anonymous):

since area is 600 put length as x and width as \[\frac{600}{x}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

then cost is \[8x+4x+4\times \frac{600}{x}+4\times \frac{600}{x}\] \[12x+\frac{4800}{x}\]